# Shamier Settle, Matthew Shapiro, Emerita Torres, and Anand Vidyashankar for discussion and feedback on the work described

### v5

@@ 4. Model Validation
We make two assumptions in Section 3. The main assumption is that the status and spatial distribution of the respondents is representative. This assumption is necessary to ensure the ratio estimator is consistent. A secondary assumption is that the spatial distribution of vendors is well described by a family of inhomogeneous Poisson processes. This assumption is necessary to derive the standard errors.
In this section, we relax both assumptions. Since neither can be relaxed from the data alone, we consider additional information. In Section 4.1, we examine a generalization of the ratio estimator in which sampling weights adjust for differential nonresponse. In Section 4.2, we model the arrival of street vendors as a pure birth process, and we derive a more general formula for the standard errors in which the amount of extra-Poisson variation can be determined by the number of vendors that cluster within markets.
@@ 4.1 Representativeness
The main assumption in Section 3 is that the status and spatial distribution of the respondents is representative. If this assumption holds, then the ratio estimator is consistent for a wider class of point processes in which Campbell’s Theorem applies, not just the Poisson process. See Daley and Vere-Jones (2007) for a general statement of Campbell’s Theorem. If this assumption is violated, then $\hat{\Lambda}_{0}(B)$ may be inconsistent. For example, if respondents are more likely to come from locations in which a higher proportion of vendors have permits, then the ratio $n_{0}(B)\,/\,n_{1}(B)$ would not be representative, and $\hat{\Lambda}_{0}(B)$ would likely underestimate $\Lambda_{0}(B)$, even in large samples.
We think the representativeness assumption is reasonable because the Street Vendor Project canvassed New York City to distribute a large amount of relief aid as discussed in Section 2. Nevertheless, to evaluate how the estimates might change when this assumption is violated, we generalize the estimator in Section 3.2 to accommodate spatially varying sampling weights. The generalization is useful for determining the sensitivity of our estimates to differential nonresponse as we demonstrate in Section 5.1.
@@
$\hat{\Lambda}_{0}^{w}(A)=\dfrac{N_{1}(A)\,n^{w}_{0}(A)}{n^{w}_{1}(A)}\qquad\text{and}\qquad\hat{p}^{w}=\frac{n^{w}_{1}(A)}{N_{1}(A)}$
Under Campbell’s Theorem, the means $\mathbb{E}[n^{w}_{0}(A)]$ and $\mathbb{E}[n^{w}_{1}(A)|N_{1}(A)]$ are
[…]
We assume the weight functions $w_{0}$ and $w_{1}$ can be chosen in such a way that the expected value is inversely proportional with the probability of a response. That is,
$p=\mathbb{E}[w_{0}(x)]\,p_{0}(x)=\mathbb{E}[w_{1}(x)]\,p_{1}(x)$
@@
$\mathbb{E}\left[n^{w}_{0}(A)\right]=p\,\Lambda_{0}(A)\qquad\text{and}\qquad\mathbb{E}\left[n^{w}_{1}(A_{i})\,|\,N_{1}(A)\right]=p\,N_{1}(A)$
and that $\hat{\Lambda}_{0}^{w}(A)$ is approximately normal with mean $\Lambda_{0}(A)$ and standard error
@@
$\hat{\mu}_{k}^{2}(A)=\sum_{x\,\in\,\pi_{k}\cap A}\mathbb{E}\bigl{[}w_{k}^{2}(x)\bigr{]}\qquad\text{and}\qquad\hat{\sigma}_{kl}(A,A)=\sum_{\begin{subarray}{c}x\,\in\,\pi_{k}\cap A\\[1.39998pt] y\,\in\,\pi_{l}\cap A\\[1.39998pt] x\neq y\end{subarray}}\mathbb{C}\bigl{(}w_{k}(x),\,w_{l}(y)\bigr{)}$
See Section 7.1.4 for details including derivations for the subregion and subtotal estimators.
@@
$\hat{\Lambda}_{0}^{w}(A)=\frac{N_{1}(A)\ n^{w}_{0}(A)}{n^{w}_{1}(A)}=\hat{\Lambda}_{0}(A)\,\frac{n^{w}_{0}(A)\,/\,n_{0}(A)}{n^{w}_{1}(A)\,/\,n_{1}(A)}$
It follows that the relative bias depends only on the ratio of the average weights $n^{w}_{0}(A)\,/\,n_{0}(A)$ and $n^{w}_{1}(A)\,/\,n_{1}(A)$. For example, if the average permitted vendor is twice as likely to respond as the average unpermitted vendors, $n^{w}_{0}(A)\,/\,n_{0}(A)$ is twice the size of $n^{w}_{1}(A)\,/\,n_{1}(A)$, and $\hat{\Lambda}_{0}(A)$ is twice the size of $\hat{\Lambda}_{0}^{w}(A)$.
@@
$\hat{\tau}^{w}(A)=\frac{N_{1}(A)\left(n^{w}_{0}(A)+n^{w}_{1}(A)\right)}{n^{w}_{1}(A)}=\hat{\tau}(A)\,\frac{\left(n^{w}_{0}(A)+n^{w}_{1}(A)\right)\,/\,\left(n_{0}(A)+n_{1}(A)\right)}{n^{w}_{1}(A)\,/\,n_{1}(A)}$
The same is true for subregion and subtotal estimators, see Section 7.1.4.
@@ 4.2 Overdispersion
We think the Poisson assumption is reasonable because a wide class of data generating processes are well-approximated by the Poisson process under the law of rare events as discussed in Section 3.2. Nevertheless, to evaluate how the standard errors might change when this assumption is violated, we consider a general model that describes how street vendors arrived at their present locations. Specifically, we consider two families of pure birth processes: one with a constant arrival rate (or “birth rate”) in which the Poisson assumption holds and one with a linear arrival rate in which the Poisson assumption does not hold.
Suppose vendors serve customers who congregate at fixed locations, which we call markets. Let $M(B)$ denote the number of markets in any subregion $B\subseteq A\subset\mathbb{R}^{2}$. If every market is in a competitive equilibrium such that the expected profit for a new vendor is the same at each market, then the time between vendor arrivals at each market may be well-approximated by an exponential distribution with a constant arrival rate $\theta_{i}$. Further suppose each market starts with one vendor at $t=0$. Then at the time the survey was conducted ($t=1$), the number of vendors, $N_{i}(B)$, would follow a Poisson distribution with mean $\Lambda_{i}(B)=\theta_{i}\,M(B)$, and the results of Section 3.2 hold.
Now suppose instead that larger markets grow faster than smaller markets, an empirical phenomenon known as preferential attachment. Then the time between vendor arrivals may be better approximated by an exponential distribution with a linear arrival rate, or Yule process. That is, let $N^{t}_{i}(B)$ denote the number of vendors in subregion $B$ at time $t$ so that the time between vendor arrivals at each market follows an exponential distribution with rate $\theta_{i}\,N^{t}_{i}(B)$. If we assume further that each market started with one vendor at $t=0$, then the number of vendors at the time the survey was conducted ($t=1$), $N_{i}(B)$, would follow a negative binomial distribution with mean $\Lambda_{i}(B)=\exp(\theta_{i})\,M(B)$ and variance $\Lambda_{i}(B)\bigr{(}\Lambda_{i}(B)-M(B)\bigr{)}\,/\,M(B)$. i.e.,
@@
where we parameterize the negative binomial distribution by its mean and variance for comparison with Section 3.2. In this case, $N_{i}(B)$ exhibits extra-Poisson variation, and the results of Section 3.2 do not hold. See Ross (2014, chap. 6) for details.
To derive the standard errors, suppose a portion of the markets, $p$, were selected for the survey at random so that
$n_{i}(B)\sim\ \text{Negative Binomial}\left(p\,\Lambda_{i}(B),\ p\,\Lambda_{i}(B)\,\frac{\Lambda_{i}(B)-M(B)}{M(B)}\right)$
@@
$n_{1}(B)\,|\,N_{1}(A)\sim\ \text{Negative Hypergeometric}\bigl{(}N_{1}(A)-1,\ M(A)-1,\ p\,q(B)\,M(A)\bigl{)}$
@@
so that under increasing domain asymptotics, the ratio estimator
@@
$\hat{u}_{0}(B)=\cfrac{\dfrac{N_{1}(A)\ n_{0}(B)}{n_{1}(A)}-M(B)}{M(B)}\qquad\text{and}\qquad u_{1}(A)=\dfrac{N_{1}(A)-M(A)}{M(A)+1}$
See Section 7.1.5 for details including derivations for the subregion and subtotal estimators. Note that our parameterization makes use of the relationship $q(B)=M(B)\,/\,M(A)$.
The standard error derived in this section for $\hat{\Lambda}_{0}(A)$ is identical to the standard error for $\hat{\Lambda}_{0}(A)$ from Section 3.2 except for the weights $u_{0}$ and $u_{1}$. These weights are the result of the extra-Poisson variation induced by preferential attachment—that is, the linear “birth” rate in which the arrival of new vendors is proportional to the number of vendors in the market. For example, if there are five vendors in each market on average, then $u_{0}$ and $u_{1}$ are approximately 4 and the standard error of the ratio estimator derived in this section is twice the size of the standard error derived in Section 3.2. The same is true for subregion and subtotal estimators, see Section 7.1.5.
@@ 5. Data Analysis
We apply the results presented in Section 3 and 4 to the data described in Section 2. In Section 5.1, we use the ratio estimator described in Section 3.1 and 3.2 to estimate the number of street vendors in New York City. In Section 5.2, we use the extensions described in Section 4.1 and 4.2 to deteremine the sensitivity of the estimate to the underlying assumptions. Sensitivity is assessed using additional data from the NYC Office of Administrative Trials and Hearings (OATH) (accessed 2024-04-03) and the American Community Survey from the U.S. Census Bureau (accessed 2022-12-22).
@@ 5.1 Estimation
According to the survey data described in Section 2.2, $n_{1}(A)=349$ mobile food vendor respondents indicated they had a permit, and $n_{0}(A)=1,051$ indicated they did not. Since there are $N_{1}(A)=5,100$ permits available, it follows that the number of vendors without permits is 15,3505,100×1,051/ 349formulae-sequence153505100105134915,350$\approx$ 5,100$\times$ 1,051\,/\,34915 , 350 ≈ 5 , 100 × 1 , 051 / 349. Combined with the $5,100$ permitted vendors, the total number of mobile food vendors is approximately $20,500$.
By similar reasoning—i.e., exploiting the fact that the number of general merchandise licenses is limited for non-veterans—the number of non-veteran merchandise vendors is estimated to be approximately $1,400$. When combined with the estimated $1,000$ veteran merchandise vendors (see Section 2.1) and $20,500$ food vendors, the total number of street vendors is approximately $23,000$.
Estimates for select subregions are listed in Table 1. We find that approximately one quarter of street vendors are located in the Manhattan neighborhoods of Chelsea, Clinton, and Lower Manhattan. Another quarter is located in West and North Queens. We also report the margin of error (i.e., two standard errors, half of the width of a 95% confidence interval).
These estimates can be checked by comparing them to a second, independent assessment conducted by the Street Vendor Project in the Bronx on May 13 and 15, 2022, in which the number of street vendors was documented through on-the-ground observations. The independent assessment identified 188 total street vendors near Fordham Road (Fordham Road BID and Street Vendor Project 2024). According to the survey data, 17 respondents indicated that they operate near Fordham Road, yielding an estimate of 166 street vendors total with a standard error of 54. We conclude this estimate is consistent with the 188 street vendors counted independently by Fordham Road BID and Street Vendor Project (2024).
{ThreePartTable}{TableNotes}
@@
These maps compare the spatial distribution of respondents (left) to the distribution of vendors suggested by administrative data (New York City Office of Administrative Trials and Hearings (OATH), top right) and federal data (American Community Survey (ACS), bottom right).
@@ 5.2 Model Validation
Nevertheless, it is possible that canvassing operations systematically missed locations. We thus compare the spatial distribution of vending locations reported by respondents to the spatial distribution of vendors who allegedly violated street vending laws during the year 2021. We consider the location of an individual’s first violation in 2021, reported in administrative data from the NYC Office of Administrative Trials and Hearings (OATH) (accessed 2024-04-03). We also compare the residencies reported by the respondents to the residencies of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers (Standard Occupational Classification 41-9091) estimated in the 2017-2021 American Community Survey (ACS) Public Use Micro Sample from the U.S. Census Bureau (accessed 2022-12-22). These comparisons are visualized in Figure 1. We find the spatial distribution of vending locations largely agree (top panels) as do the spatial distribution of the residencies (bottom panels). This suggests that, if there were a region of street vendors missed by the SVP, it was also missed by law enforcement and the ACS.
@@
We can also combine the data from OATH and ACS with the results from Section 4.1 to assess the sensitivity of the estimates in Section 5.1 to differential nonresponse, as we now demonstrate. While the difference between these new estimates and the estimates provided in Section 5.1 are not statistically significant at the 95% level, we believe they provide insight into how the estimates in Section 5.1 might theoretically err.
@@
First, we consider the scenario in which the probability a vendor responds is inversely proportional to the average number of vendors cited for violating street vending laws in the surrounding ZIP Code each day, as reported in the OATH data for the year 2021. This might be the case if vendors in high enforcement areas are reluctant to interact with survey takers. The weights are the inverse probabilities—the average number of enforcement actions taken against vendors in the surrounding ZIP Code. We find the weighted counts for food vendors are $n^{w}_{0}(A)=25.7$ and $n^{w}_{1}(A)=5.24$. This suggests there are $\left(n_{0}^{w}(A)+n_{1}^{w}(A)\right)\,/\,\left(n_{0}(A)+n_{1}(A)\right)\,/\left(n_{1}^{w}(A)\,/\,n_{1}(A)\right)=47\%$ more food vendors than estimated in Section 5.1. A similar argument for merchandise vendors suggests there are 9% more merchandise vendors than estimated in Section 5.1. Put together, the weights suggest the number of street vendors in New York City is 40% higher than estimated in Section 5.2.
@@
The second scenario we consider is one in which the probability a vendor responds in proportional to the average number of vendors cited for violating street vending laws in the surrounding ZIP Code. This might be the case if vendors in high enforcement areas are easier to identify and thus easier to survey. Under this scenario, there are 8% fewer food vendors than estimated in Section 5.1 and 9% fewer merchandise vendors, suggesting the number of street vendors in New York City could be 8% lower than estimated in Section 5.2.
@@
The final two scenarios are identical to the first two scenarios, except we use the number of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers in the surrounding Public Use Microdata Area as estimated from the 2017-2021 U.S. Census Bureau (accessed 2022-12-22). The result is less extreme than using the average number of vendors cited. If the probability of response is inversely proportional to this estimate, there are 33% more vendors New York City than estimated in Section 5.2. If the probability of response is proportional to this estimate, there are 5% fewer vendors. Combining these four scenarios suggests a conservative range of between 21,000 and 32,200 street vendors operate in New York City.
@@
The second assumption is that the spatial distribution of vendors is well described by a family of inhomogeneous spatial Poisson processes. We argue that this assumption is reasonable because a wide class of data generating processes are well-approximated by the Poisson process under the law of rare events as discussed in Section 3.2.
@@
It is possible, however, that vendors cluster within markets as described in Section 4.2. The weights $u_{0}$ and $u_{1}$ in Section 4.2 correspond to the relative number of vendors without permits and with permits per market respectively. These weights cannot be estimated from the survey data alone. However, we believe that $\Lambda_{0}(A)$ and $N_{1}(A)$ are no more than 10 times larger than $M(A)$. It follows that $u_{0}$ and $u_{1}$ are less than 9, and thus the standard errors from Section 3.2 and the margins of error from Table 1 are at most 3 times larger if the Poisson assumption is violated and the Yule process holds. We suspect this bound is conservative, however, and the typical market contains only one or two vendors.

### v6

@@ 4. Model Validation
We make two assumptions in Section 3. The main assumption is that the status and spatial distribution of the respondents is representative. This assumption is sufficient to ensure the ratio estimator is consistent. A secondary assumption is that the spatial distribution of vendors is well approximated by a family of inhomogeneous Poisson processes. This assumption is sufficient to derive the standard errors analytically.
In this section, we relax both assumptions. Since neither can be relaxed from the data alone, we consider additional information. In Section 4.1, we examine a generalization of the ratio estimator in which sampling weights adjust for differential nonresponse. In Section 4.2, we model the spatial distribution of respondents using an inhomogeneous negative binomial process, and we derive a more general formula for the standard errors in which extra-Poisson variation or “overdispersion” can be explained by the number of vendors that cluster within markets. The details are outlined in Sections 7.4 and 7.5.
@@ 4.1 Representativeness
The main assumption in Section 3 is that the status and spatial distribution of the respondents is representative. If this assumption holds, then the ratio estimator is consistent for a wider class of point processes in which Campbell’s Theorem applies, not just the Poisson process. See Daley and Vere-Jones (2007) for a general statement of Campbell’s Theorem. If this assumption is violated, then $\hat{\Lambda}_{0}(B)$ may be inconsistent. For example, if respondents are more likely to come from locations in which a higher proportion of vendors have permits, then the ratio $n_{0}(B)\,/\,n_{1}(B)$ would not be representative and $\hat{\Lambda}_{0}(B)$ would likely underestimate $\Lambda_{0}(B)$, even in large samples.
We think the representativeness assumption is reasonable because the Street Vendor Project canvassed New York City to distribute a large amount of relief aid as discussed in Section 2. Nevertheless, to evaluate how the estimates might change when this assumption is violated, we generalize the estimator in Section 3.2 to accommodate spatially varying sampling weights. The generalization allows us to determine the sensitivity of our estimates to differential nonresponse as we demonstrate in Section 5.1.
@@
$\hat{\Lambda}_{0}^{w}(A)=\dfrac{N_{1}(A)\,n^{w}_{0}(A)}{n^{w}_{1}(A)}\qquad\text{and}\qquad\hat{p}^{w}=\frac{n^{w}_{1}(A)}{N_{1}(A)}.$
Under Campbell’s Theorem, the expected values $\mathbb{E}[n^{w}_{0}(A)]$ and $\mathbb{E}[n^{w}_{1}(A)\,|\,N_{1}(A)]$ are
$\mathbb{E}\left[n^{w}_{0}(A)\right]=\int_{A}\mathbb{E}\bigl[w_{0}(x)\bigr]\,p_{0}(x)\,\lambda_{0}(x)\,dx\qquad\text{and}\qquad\mathbb{E}\left[n^{w}_{1}(A)\,|\,N_{1}(A)\right]=\frac{N_{1}(A)}{\Lambda_{1}(A)}\,\int_{A}\mathbb{E}\bigl[w_{1}(x)\bigr]\,p_{1}(x)\,\lambda_{1}(x)\,dx.$
We assume the weight functions $w_{0}$ and $w_{1}$ can be chosen in such a way that the expected values are inversely proportional to the probability of a response. That is,
$p=\mathbb{E}[w_{0}(x)]\,p_{0}(x)=\mathbb{E}[w_{1}(x)]\,p_{1}(x).$
This choice is similar in spirit to Horvitz and Thompson (1952) in that we seek to correct for the influence of $p_{0}(x)$ and $p_{1}(x)$. Note that we only require the average weight to be inversely proportional to the probability of a response.
@@
$\mathbb{E}\left[n^{w}_{0}(A)\right]=p\,\Lambda_{0}(A)\qquad\text{and}\qquad\mathbb{E}\left[n^{w}_{1}(A)\,|\,N_{1}(A)\right]=p\,N_{1}(A)$
and that $\hat{\Lambda}_{0}^{w}(A)$ is asymptotically normal with mean $\Lambda_{0}(A)$ and standard error
@@
$\hat{\mu}_{k}^{2}(A)=\sum_{x\,\in\,\pi_{k}\cap A}\mathbb{E}\bigl[w_{k}^{2}(x)\bigr]\qquad\text{and}\qquad\hat{\sigma}_{kl}(A,A)=\sum_{\begin{subarray}{c}x\,\in\,\pi_{k}\cap A\\[1.63885pt] y\,\in\,\pi_{l}\cap A\\[1.63885pt] x\neq y\end{subarray}}\mathbb{C}\bigl(w_{k}(x),\,w_{l}(y)\bigr).$
See Section 7.4 for details including derivations for the subregion and subtotal estimators.
@@
$\hat{\Lambda}_{0}^{w}(A)=\frac{N_{1}(A)\ n^{w}_{0}(A)}{n^{w}_{1}(A)}=\hat{\Lambda}_{0}(A)\,\frac{n^{w}_{0}(A)\,/\,n_{0}(A)}{n^{w}_{1}(A)\,/\,n_{1}(A)}.$
It follows that the relative bias depends only on the ratio of the average weights $n^{w}_{0}(A)\,/\,n_{0}(A)$ and $n^{w}_{1}(A)\,/\,n_{1}(A)$. For example, if the average permitted vendor is twice as likely to respond as the average unpermitted vendors, $n^{w}_{0}(A)\,/\,n_{0}(A)$ is twice the size of $n^{w}_{1}(A)\,/\,n_{1}(A)$ and $\hat{\Lambda}_{0}(A)$ is twice the size of $\hat{\Lambda}_{0}^{w}(A)$.
@@
$\hat{\tau}^{w}(A)=\frac{N_{1}(A)\left(n^{w}_{0}(A)+n^{w}_{1}(A)\right)}{n^{w}_{1}(A)}=\hat{\tau}(A)\,\frac{\left(n^{w}_{0}(A)+n^{w}_{1}(A)\right)\,/\,\left(n_{0}(A)+n_{1}(A)\right)}{n^{w}_{1}(A)\,/\,n_{1}(A)}.$
The same is true for the subregion and subtotal estimators, see Section 7.4.
@@ 4.2 Overdispersion
We think the Poisson assumption is reasonable because a wide class of data generating processes are well-approximated by a Poisson process under the law of rare events as discussed in Section 3.2. Nevertheless, to evaluate how the standard errors might change when this assumption is violated, we consider an alternative justification based on the theory of birth processes. As in Section 3.2, we only sketch the justification. See Ross (2014, chap. 6) for an extended discussion of the theory of birth processes.
This second justification asks how street vendors arrived at their present locations, and we consider two possible explanations. We first suppose that vendors serve customers who congregate at fixed locations, which we call markets. Let $M(B)$ denote the number of markets in any subregion $B\subseteq A\subset\mathbb{R}^{2}$. We assume markets grow over time as vendors arrive to serve customers. If vendor arrivals are well approximated by an exponential distribution with a constant arrival rate or “birth rate”, then the number of vendors at the time of the survey, $N_{i}(B)$, would follow a Poisson distribution as assumed in Section 3.2.
Suppose instead that larger markets grow faster than smaller markets, an empirical phenomenon known as preferential attachment. Then the time between vendor arrivals may be better approximated by an exponential distribution with a linear arrival rate or Yule process. In this case, the number of vendors at the time of the survey, $N_{i}(B)$, would follow a negative binomial distribution with mean $\Lambda_{i}(B)$ and variance $\Lambda_{i}(B)\bigr(\Lambda_{i}(B)-M(B)\bigr)\,/\,M(B)$. i.e.,
@@
where we parameterize the negative binomial distribution by its mean and variance for comparison with Section 3.2. In this case, $N_{i}(B)$ exhibits extra-Poisson variation or “overdispersion” when many vendors cluster within a market and thus the results of Section 3.2 may not hold. Note that we define the negative binomial as the total number of trials required to get a predetermined number of successes when sampling with replacement.
To derive the standard errors, suppose a portion of the markets, $p$, were selected for the survey so that
$n_{i}(B)\sim\ \text{Negative Binomial}\left(p\,\Lambda_{i}(B),\ p\,\Lambda_{i}(B)\,\frac{\Lambda_{i}(B)-M(B)}{M(B)}\right).$
@@
$n_{1}(B)\,|\,N_{1}(A)\sim\ \text{Negative Hypergeometric}\bigl(N_{1}(A)-1,\ M(A)-1,\ p\,q(B)\,M(A)\bigl).$
@@
so that the ratio estimator
@@
$\hat{u}_{0}(B)=\cfrac{\dfrac{N_{1}(A)\ n_{0}(B)}{n_{1}(A)}-M(B)}{M(B)}\qquad\text{and}\qquad u_{1}(A)=\dfrac{N_{1}(A)-M(A)}{M(A)+1}.$
See Section 7.5 for details including derivations for the subregion and subtotal estimators. Note that our parameterization makes use of the relationship q(B)=M(B)/M(A)q(B)=M(B)\,/\,M(A).
The standard error derived in this section for $\hat{\Lambda}_{0}(A)$ is identical to the standard error for $\hat{\Lambda}_{0}(A)$ from Section 3.2 except for the weights $u_{0}$ and $u_{1}$. These weights can be explained by preferential attachment—the linear “birth” rate in which the number of new vendors is proportional to the number of vendors already in the market. If there are five vendors in each market on average, then $u_{0}$ and $u_{1}$ are approximately 4 and the standard error of the ratio estimator derived in this section is twice the size of the standard error derived in Section 3.2. The same is true for subregion and subtotal estimators, see Section 7.5.
@@ 5. Data Analysis
We apply the methodology developed in Section 3 and 4 to the data described in Section 2. In Section 5.1, we use the ratio estimator from Section 3.1 and 3.2 to estimate the number of street vendors in New York City. In Section 5.2, we use the extensions from Section 4.1 and 4.2 to determine the sensitivity of the estimate to the underlying assumptions. Sensitivity is assessed using additional data from the NYC Office of Administrative Trials and Hearings (OATH) (accessed 2024-04-03) and the American Community Survey from the U.S. Census Bureau (accessed 2022-12-22). We also compare our estimates to those obtained using a hierarchical model.
@@ 5.1 Estimation
According to the survey data described in Section 2.2, $n_{1}(A)=349$ food vendor respondents indicated they had a permit and $n_{0}(A)=1,051$ indicated they did not. Since there are $N_{1}(A)=5,100$ permits available, it follows that the number of vendors without permits is $15,350\approx 5,100\times 1,051\,/\,349$. Combined with the $5,100$ permitted vendors, the total number of mobile food vendors is approximately $20,500$.
By similar reasoning—i.e., exploiting the fact that the number of merchandise licenses is limited for non-veterans—the number of non-veteran merchandise vendors is estimated to be approximately $1,400$. When combined with the estimated $1,000$ veteran merchandise vendors (see Section 2.1) and $20,500$ food vendors, the total number of street vendors is approximately $23,000$.
Estimates for select subregions are listed in Table 1. We find that approximately one quarter of street vendors are located in the Manhattan neighborhoods of Chelsea, Clinton, and Lower Manhattan. Another quarter is located in West and North Queens. We also report the margin of error (i.e., two standard errors, half of the width of a 95% confidence interval). The margin of error is calculated using the asymptotic approximations from Section 3.2. The subregions are chosen so that the estimated number of vendors in each region is sufficiently large that the asymptotic approximations are expected to be accurate.
We check our estimates by comparing them to a second, independent assessment conducted by the Street Vendor Project in the Bronx on May 13 and 15, 2022, in which the number of street vendors was documented through on-the-ground observations. The independent assessment identified 188 total street vendors near Fordham Road (Fordham Road BID and Street Vendor Project 2024). According to the survey data, 17 respondents indicated that they operate near Fordham Road, yielding an estimate of 166 street vendors total with a standard error of 54. We conclude this estimate is consistent with the 188 street vendors counted independently by Fordham Road BID and Street Vendor Project (2024).
@@
These maps compare the spatial distribution of respondents (left) to the distribution of vendors suggested by administrative data (New York City Office of Administrative Trials and Hearings (OATH), top right) and federal data (American Community Survey (ACS), bottom right). The top two maps show the log number of vendors by the ZIP Code of their vending location, and the bottom two maps show the log number of vendors by the PUMA (Public Use Microdata Area) of their residency. The top two maps show a similar pattern in which vendors concentrate in North and West Queens and in the Chelsea and Clinton neighborhoods of Manhattan. The bottom two maps show a similar pattern in which residencies are more uniformly distributed in North and West Queens, the South Bronx, and Upper Manhattan.
@@ 5.2 Model Validation
Nevertheless, it is possible that canvassing operations systematically missed locations. We thus compare the spatial distribution of vending locations reported by respondents to the spatial distribution of vendors who allegedly violated street vending laws during the year 2021. We consider the location of an individual’s first violation in 2021, reported in administrative data from the NYC Office of Administrative Trials and Hearings (OATH) (accessed 2024-04-03). We also compare the residencies reported by the respondents to the residencies of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers (Standard Occupational Classification 41-9091) estimated in the 2017-2021 American Community Survey (ACS) Public Use Micro Sample from the U.S. Census Bureau (accessed 2022-12-22). These comparisons are visualized in Figure 1. We find the spatial distribution of vending locations largely agree (top panels) as do the spatial distribution of the residencies (bottom panels) in that discrepancies are relatively minor and explainable by chance. We conclude that if there were a region of street vendors missed by the SVP, it was also missed by law enforcement and the ACS.
@@
We can also combine the data from OATH and ACS with the results from Section 4.1 to assess the sensitivity of the estimates in Section 5.1 to differential nonresponse as we now demonstrate. While the difference between these new estimates and the estimates provided in Section 5.1 are not statistically significant at the 95% level, we believe they provide insight into how the estimates in Section 5.1 might theoretically err.
@@
First, we consider the scenario in which the probability a vendor responds is inversely proportional to the average number of vendors cited for violating street vending laws in the surrounding ZIP Code each day. This might be the case if vendors in high enforcement areas are reluctant to interact with survey takers. The weights are the inverse probabilities—the average number of enforcement actions taken against vendors in the surrounding ZIP Code. We find the weighted counts for food vendors are $n^{w}_{0}(A)=25.7$ and $n^{w}_{1}(A)=5.24$. This suggests there are
$\dfrac{\left(n_{0}^{w}(A)+n_{1}^{w}(A)\right)\,/\,\left(n_{0}(A)+n_{1}(A)\right)}{n_{1}^{w}(A)\,/\,n_{1}(A)}=47\%$
more food vendors than estimated in Section 5.1. A similar argument for merchandise vendors suggests there are 9% more merchandise vendors than estimated in Section 5.1. Put together, the weights suggest the number of street vendors in New York City is 40% higher than estimated in Section 5.2.
@@
The second scenario we consider is one in which the probability a vendor responds is proportional to the average number of vendors cited for violating street vending laws in the surrounding ZIP Code. This might be the case if vendors in high enforcement areas are easier to identify and thus easier to survey. In this scenario, there are 8% fewer food vendors than estimated in Section 5.1 and 9% fewer merchandise vendors, suggesting the number of street vendors in New York City could be 8% lower than estimated in Section 5.2.
@@
The final two scenarios are identical to the first two, except we use the number of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers in the surrounding Public Use Microdata Area as estimated from the 2017-2021 U.S. Census Bureau (accessed 2022-12-22) (as indicated by the POWPUMA variable). The result is less extreme than using the average number of vendors cited. If the probability of response is inversely proportional, there are 33% more vendors in New York City than estimated in Section 5.2. If the probability of response is proportional, there are 5% fewer vendors. Combining these four scenarios suggests a conservative range of between 21,000 and 32,200 street vendors operate in New York City. The lower and upper limit of this range are estimates. Using the standard deviations in Section 4.1, a conservative 95% confidence interval is 17,200 to 41,800.
@@
The second assumption is that the spatial distribution of vendors is well approximated by a family of inhomogeneous Poisson processes. We argue this assumption is reasonable because a wide class of data generating processes are well-approximated by a Poisson process under the law of rare events as discussed in Section 3.2. The fact that many street vendors operate in a few Zip Codes can be explained by socioeconomic factors unique to those Zip Codes and does not necessarily suggest a more complicated model.
@@
Nevertheless, it is possible that vendors cluster within markets according to a birth process as described in Section 4.2. In that case, the spatial distribution of vendors may be better approximated by a family of inhomogeneous negative binomial processes where the weights $u_{0}$ and $u_{1}$ correspond to the relative number of vendors without permits and with permits per market respectively. These weights cannot be estimated from the survey data alone. However, we believe that if there is clustering as described in Section 4.2, $\Lambda_{0}(A)$ and $N_{1}(A)$ are no more than 10 times larger than $M(A)$. It follows that $u_{0}$ and $u_{1}$ are less than 9, and thus the standard errors from Section 3.2 and the margins of error from Table 1 are at most 3 times larger. We suspect this bound is conservative, however, and the typical market contains only one or two vendors.
We conclude this section by comparing our results to a hierarchical approach in which both the expected number of vendors and the probability a vendor responds to the survey varies by subregion. The hierarchical approach provides an alternative framework for constructing estimates and standard errors, and comparison with the proposed approach allows us to double check the sensitivity of our estimates and standard errors to our underlying assumptions. The hierarchical approach still assumes the independent Poisson processes and their thinned counterparts as described in Section 3.2. However, we depart from Section 3.2 by allowing $p(B_{i})$, $\Lambda_{0}(B_{i})$, and $\Lambda_{1}(B_{i})$ to vary according to the model
$\text{logit}\,p(B_{i})\sim$ $\ \text{Normal}(\mu_{p},\,\sigma_{p}^{2})$
$\text{log}\,\Lambda_{0}(B_{i})\sim$ $\ \text{Normal}(\mu_{0},\,\sigma_{0}^{2})$
$\text{log}\,\Lambda_{1}(B_{i})\sim$ $\ \text{Normal}(\mu_{1},\,\sigma_{1}^{2}).$
Note that the hierarchical approach also departs from Sections 4.1 and 4.2 because $\Lambda_{0}(B_{i})$ and $\Lambda_{1}(B_{i})$ now share a common distribution across subregions so that the hyperparameters $\mu_{p}$ and $\sigma_{p}^{2}$ are identified.
We use the data described in Section 2 and the statistical software Stan to sample from the posterior distribution of the hierarchical model. We then use the posterior draws to obtain estimates and uncertainty intervals. The posterior mean of the total number of vendors is 21,000 with a 95% credible interval of 18,900 to 23,900, which is consistent with the proposed approach. A similar hierarchical model using the negative binomial process introduced in Section 4.2 produces a similar interval. See Section 7.6 for details including the Stan code used.